OpenStudy (anonymous):

The sum of two numbers is 77. The greater number is 5 more than the smaller number. Which equation can be used to solve for the smaller number?

OpenStudy (anonymous):

x(x − 5) = 77 x(x + 5) = 77 x − (x + 5) = 77 x + (x + 5) = 77

OpenStudy (jackthegreatest):

X + x + 5 = 75

OpenStudy (jackthegreatest):

Wait I meant 77

OpenStudy (jackthegreatest):

So d

OpenStudy (anonymous):

its D

OpenStudy (xapproachesinfinity):

D is good

OpenStudy (xapproachesinfinity):

always try to think in algebraic way. when the say the number (not saying what is ) give it a name x or whatever letter denoting that it is what we are looking for so we have in this case two number if said the small number is x and we know the greater is 5 more than the small we are essentially saying x+5 and the sum of the two x+(x+5)=77

OpenStudy (xapproachesinfinity):

it is a matter of reading and translating the meaning of a words within the problem